OpenStudy (anonymous):

Use graph to evaluate sin(4π)

OpenStudy (anonymous):

Please help i have no idea how to do this

zepdrix (zepdrix):

Is there a graph provided with the problem?

OpenStudy (anonymous):

No it just says use a graph and thats whats confusing me

OpenStudy (anonymous):

Any idea?

zepdrix (zepdrix):

Ok let's try using this graph then. https://www.desmos.com/calculator/8j4phwyggg Here I have graphed the function \(\large y=\sin(x)\) They want to know, when \(\large x=4\pi\), where does the \(\large y\) value land? Can you tell based on this graph?

OpenStudy (anonymous):

0?

zepdrix (zepdrix):

Yah it looks like it's going right through the \(\large x-axis\) when we go through \(\large 4\pi\) right? Yes, good job! :)

OpenStudy (anonymous):

Thank you so much, ive been stuck on this problem for a while haha

zepdrix (zepdrix):

heh np c: